# Complexity Bounds for Smooth Convex Multiobjective Optimization

### v1

@@ Abstract
We study the oracle complexity of finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization with $m$ objectives. The progress metric is the Pareto stationarity gap $\mathcal{G}(x)$ (the norm of an optimal convex combination of gradients). Our contributions are fourfold. (i) For strongly convex objectives, any span first-order method (iterates lie in the span of past gradients) exhibits linear convergence no faster than $\exp(-\Theta(T/\sqrt{\kappa}))$ after $T$ oracle calls, where $\kappa$ is the condition number, implying $\Theta(\sqrt{\kappa}\log(1/\varepsilon))$ iterations; this matches classical accelerated upper bounds. (ii) For convex problems and oblivious one-step methods (a fixed scalarization with pre-scheduled step sizes), we prove a lower bound of order $1/T$ on the best gradient norm among the first $T$ iterates. (iii) Although accelerated gradient descent is outside this restricted class, it is an oblivious span method and attains the same $1/T$ upper rate on a fixed scalarization. (iv) For convex problems and general span methods with adaptive scalarizations, we establish a universal lower bound of order $1/T^{2}$ on the gradient norm of the final iterate after $T$ steps, highlighting a gap between known upper bounds and worst-case guarantees. All bounds hold on non-degenerate instances with distinct objectives and non-singleton Pareto fronts; rates are stated up to universal constants and natural problem scaling.
@@ 1 Introduction
Why scalarization does not trivialize MOO.
It may appear that multiobjective complexity reduces immediately to the single-objective case by fixing a scalarization $f_{\lambda}=\sum_{i}\lambda_{i}f_{i}$ and applying known convergence results. Indeed, all classical rates for smooth convex or strongly convex minimization (e.g., gradient descent $O(LR/\sqrt{T})$, AGD $O(LR/T)$, and linear convergence under strong convexity) apply verbatim to any fixed scalarization. However, the multiobjective goal is different: the central complexity measure is the Pareto stationarity gap $\mathcal{G}(x)$ which requires considering all convex combinations of gradients simultaneously. A point minimizing a fixed $f_{\lambda}$ need not minimize $\mathcal{G}$, and the algorithm classes we study (oblivious one-step vs. span/adaptive methods) differ in their ability to exploit information across objectives. Consequently, while scalarization provides useful upper bounds, the oracle complexity of Pareto stationarity requires a dedicated analysis, which is the focus of this paper.
Contributions.
@@
  • Universal Lower Bound for Adaptive Methods: For fully adaptive span first-order methods, we establish a universal last-iterate lower bound of $\Omega(LR/(T+1)^{2})$. We highlight the remaining gap between this and the $\mathcal{O}(1/T)$ upper bounds of optimal methods as a key open problem.
  • Robustness of Constructions: We show our lower bounds are not artifacts of degenerate problems. They hold for instances with distinct, conflicting objectives and non-singleton Pareto fronts.
Our methodology refines the classical resisting oracle framework [23] by carefully analyzing the polynomial representations of different algorithm classes. These results provide a definitive characterization of the oracle complexity for a central task in modern multiobjective optimization.
Organization.
After surveying the literature in the Related Work section, Section 3 fixes notation, introduces the Pareto stationarity gap $\mathcal{G}(\cdot)$ and its equivalence to Pareto criticality, and formalizes the oracle model and algorithm classes (span methods vs. oblivious one-step), including the Krylov/span polynomial representation. Section 4 develops our lower bounds: a tight linear-rate bound in the strongly convex case (Theorem 4.1), an $\Omega(1/T)$ min-iterate bound for oblivious one-step methods (Theorem 4.3), and a universal $\Omega(1/T^{2})$ last-iterate bound for fully adaptive span methods (Theorem 4.4), together with an explicit non-degenerate construction showing robustness of the instances (Section 4.3). Section 5 provides matching upper bounds via AGD on fixed scalarizations and compares rates against the lower bounds. Section 6 concludes with implications and open questions. The appendices contain the polynomial extremal derivations (Chebyshev and Markov) and auxiliary proofs used throughout.
@@ 2 Related Work
Oracle complexity and first-order limits.
@@
The information-based complexity program for single-objective convex optimization established sharp first-order lower bounds together with matching optimal methods [23, 24]. In [23], these lower bounds are derived by reducing the worst-case error to an extremal polynomial problem, solved exactly by (scaled) Chebyshev polynomials; see [28] for the underlying Chebyshev minimax and related inequalities. Recent analyses include Performance Estimation Problem (PEP) frameworks that deliver tight worst-case guarantees for many first-order methods [8]. Beyond the classical oracle model, a line of work formalizes oblivious and structured first-order algorithms: Arjevani and Shamir [3] introduced the oblivious framework with dimension-free lower bounds for $L$-smooth convex and $L$-smooth $\mu$-strongly convex problems; Arjevani, Shalev-Shwartz, and Shamir [1] developed the $p$-SCLI framework (covering p=1p{=}1 one-step methods) with matching bounds; and Arjevani and Shamir [2] extended dimension-free iteration lower bounds to finite-sum settings, covering variance-reduced families. Complementary results for randomized first-order algorithms appear in Woodworth and Srebro [32]. A comprehensive and recent treatment of finite-sum lower bounds is given by Han, Xie, and Zhang [18].
Multiobjective optimization algorithms.
@@
Classical foundations and broad context are covered in the monograph by Miettinen [21] and the survey of Marler and Arora [20]. For descent-type methods, the steepest–descent family beginning with [13] and including cone-ordered variants and refined convergence analyses [9, 7] provides global convergence to Pareto-stationary points. Complexity aspects for gradient descent in smooth MOO are analyzed in [11]. Universal nonmonotone line-search frameworks for non-convex MOO with convex constraints have also been proposed [27]. When progress is tracked through standard merit functions or stationarity surrogates, the resulting upper bounds are typically sublinear (e.g., $\mathcal{O}(1/k)$) in smooth convex settings [31, 16]. Second-order frameworks — Newton-type and SQP — offer globalization with fast local behavior: globally convergent Newton methods for MOO with fast (often superlinear) local rates under standard regularity are developed in [17], while SQP-based schemes for constrained problems are given in [12]. Conditional-gradient (Frank–Wolfe) methods have been developed for vector/multiobjective optimization with global convergence to weakly efficient (Pareto-stationary) solutions under cone-convexity assumptions [6]. Away-step and related variants can accelerate convergence and, under additional geometric or curvature assumptions, deliver faster (sometimes linear) rates [15, 16], with adaptive generalized conditional-gradient methods extending this line [14]. Accelerated first-order schemes tailored to the multiobjective setting continue to mature: an accelerated proximal-gradient method with $\mathcal{O}(1/k^{2})$ decrease for an appropriate merit measure is established in [30], and recent work clarifies monotonicity properties of multiobjective APG [25]. Complementary practical gradient-based schemes and stationarity criteria, such as multiple-gradient descent, are surveyed in [10], and links to multi-task learning further motivate stationarity-based goals and algorithm design [29].
Our position.
The works above emphasize algorithm design and, in the scalar case, tight lower bounds under various oracle or structural models. To our knowledge, a systematic oracle-complexity theory for Pareto stationarity in smooth MOO — distinguishing oblivious one-step methods from general span first-order methods and providing robust non-degenerate lower-bound instances — has not been previously developed. The present paper fills this gap and connects multiobjective first-order complexity to the classical scalar theory through polynomial extremal tools.
@@ 3.1 Notation
For multiobjective problems with $m\geq 2$ objectives $F(x)=(f_{1}(x),\ldots,f_{m}(x))$, the unit simplex is $\Delta_{m}:=\{\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\}$. For $\lambda\in\Delta_{m}$, the scalarization is $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ and $x^{\star}_{\lambda}\in\arg\min_{x}f_{\lambda}(x)$ denotes an arbitrary minimizer. The Pareto set is denoted by $\mathcal{P}$. We use $R:=\operatorname{dist}(x(0),\mathcal{P})$ and $R_{\lambda}:=\|x(0)-x^{\star}_{\lambda}\|$ for initialization radii. Iterates are $x(t)$ for $t\in\mathbb{N}_{0}$; $T\in\mathbb{N}$ is the number of oracle calls/iterations. The condition number is $\kappa:=L/\mu$ when strong convexity applies; we occasionally write $\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}$ in spectral formulas. For quadratics $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, we use the spectral variable $\zeta$ and real polynomials $p_{t}$ (with $\deg p_{t}\leq t$, $p_{t}(0)=1$) to express errors via $x(t)-x^{\star}=p_{t}(H)\bigl{(}x(0)-x^{\star}\bigr{)}$; Chebyshev polynomials of the first kind are denoted $T_{t}(\cdot)$. Standard Landau symbols $O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ are used; $\log$ denotes the natural logarithm. The vectors $(e_{i})_{i=1}^{m}$ are the canonical basis of $\mathbb{R}^{m}$.
@@ 3.2 Problem formulation
We consider unconstrained, continuously differentiable multiobjective problems
@@
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $L$-smooth.
@@ 3.3 Pareto optimality and first-order conditions
Equivalence of stationarity notions.
@@
    (Convex-hull stationarity) There exists $\lambda\in\Delta^{m}$ such that $\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)=0$ (equivalently, $\mathcal{G}(x)=0$).
@@
Necessary and sufficient conditions.

### v2

# Complexity Bounds for Smooth Multiobjective Optimization

@@ Abstract
We study the oracle complexity of finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization with $m$ objectives. Progress is measured by the Pareto stationarity gap $\mathcal{G}(x)$, the norm of the best convex combination of objective gradients. Our analysis relies on a non-degenerate lifting that embeds hard single-objective instances into MOO instances with distinct objectives and non-singleton Pareto fronts while preserving lower bounds on $\mathcal{G}$. We establish: (i) in the $\mu$-strongly convex case, any span first-order method has worst-case linear convergence no faster than $\exp(-\Theta(T/\sqrt{\kappa}))$ after $T$ oracle calls, yielding $\Theta(\sqrt{\kappa}\log(1/\varepsilon))$ iterations and matching accelerated upper bounds; (ii) in the convex case, an $\Omega(1/T)$ min-iterate lower bound for oblivious one-step methods and a universal last-iterate lower bound $\Omega(1/T^{2})$ for oblivious span methods via polynomial-degree arguments, and we further show this latter bound is loose (for general adaptive methods) by importing geometric lower bounds to obtain an $\Omega(1/T)$ min-iterate lower bound for general adaptive first-order methods; (iii) in the nonconvex case with $L$-Lipschitz gradients, an $\Omega(\sqrt{L}/(T+1))$-type lower bound on $\mathcal{G}$ (tight in order), implying $\Omega(1/\varepsilon^{2})$ iterations to reach $\mathcal{G}(x)\leq\varepsilon$ up to natural scaling.
@@ 1 Introduction
Why scalarization does not trivialize MOO
It may appear that multiobjective complexity reduces immediately to the single-objective case by fixing a scalarization $f_{\lambda}=\sum_{i}\lambda_{i}f_{i}$ and applying known convergence results. Indeed, all classical rates for smooth convex or strongly convex minimization (e.g., gradient descent $O(LR/\sqrt{T})$, AGD $O(LR/T)$, and linear convergence under strong convexity) apply verbatim to any fixed scalarization. However, the multiobjective goal is different: the central complexity measure is the Pareto stationarity gap $\mathcal{G}(x)$ which requires considering all convex combinations of gradients simultaneously. A point minimizing a fixed $f_{\lambda}$ need not minimize $\mathcal{G}$, and the algorithm classes we study (oblivious one-step vs. oblivious span vs. general adaptive first-order methods) differ in their ability to exploit information across objectives. Consequently, while scalarization provides useful upper bounds, the oracle complexity of Pareto stationarity requires a dedicated analysis, which is the focus of this paper.
Contributions
@@
  • Robust non-degenerate lifting. We show our lower bounds are not artifacts of degenerate problems. We construct MOO instances with distinct objectives and non-singleton Pareto fronts that preserve lower bounds on $\mathcal{G}$ from embedded scalar instances.
@@
  • Polynomial-degree lower bound for oblivious span methods and geometric lower bound for adaptive methods: Within the oblivious span model, we establish a universal last-iterate polynomial-degree lower bound $\Omega(1/T^{2})$ via Markov-type extremal inequalities. For general deterministic adaptive first-order methods, we prove a stronger geometric (zero-chain) $\Omega(1/T)$ min-iterate lower bound under a bounded scalarization gap, obtained by lifting the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [7].
  • Nonconvex extension with Lipschitz gradients. Using the same lifting on the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [6], we obtain a nonconvex MOO lower bound under $L$-Lipschitz gradients of $\Omega(\sqrt{L\Delta/(T+1)})$ (equivalently $\Omega(L\Delta/\varepsilon^{2})$ iterations) to reach $\mathcal{G}(x)\leq\varepsilon$.
Our methodology combines classical polynomial-based lower-bound techniques in the spirit of the resisting-oracle framework [27] with a non-degenerate lifting reduction that transfers sharp single-objective stationary-point lower bounds (e.g., [7]) to multiobjective Pareto stationarity. This yields a unified oracle-complexity perspective for smooth MOO and clarifies how algorithmic restrictions (oblivious one-step vs. oblivious span vs. general adaptive first-order methods) and iterate selection (min vs. last) shape the achievable rates.
Organization
Section 2 reviews related work. Section 3 fixes notation, defines $\mathcal{G}$, and formalizes oracle models and algorithm classes. Section 4 presents the lifting construction, then derives strongly convex, convex, and nonconvex lower bounds by combining polynomial extremal arguments and geometric stationary-point lower bounds. Section 5 provides matching upper bounds via AGD on fixed scalarizations and compares rates against the lower bounds. Section 6 concludes with implications and open questions. The appendices contain the polynomial extremal derivations (Chebyshev and Markov) used to analyze span and oblivious methods on hard quadratic instances.
@@ 2 Related Work
Oracle complexity and first-order limits
@@
The information-based complexity program for single-objective convex optimization established sharp first-order lower bounds together with matching optimal methods [27, 28]. In [27], these lower bounds are derived by reducing the worst-case error to an extremal polynomial problem, solved exactly by (scaled) Chebyshev polynomials; see [32] for the underlying Chebyshev minimax and related inequalities. Recent analyses include Performance Estimation Problem (PEP) frameworks that deliver tight worst-case guarantees for many first-order methods [10]. Beyond the classical oracle model, a line of work formalizes oblivious and structured first-order algorithms: Arjevani and Shamir [3] introduced the oblivious framework with dimension-free lower bounds for $L$-smooth convex and $L$-smooth $\mu$-strongly convex problems; Arjevani, Shalev-Shwartz, and Shamir [1] developed the $p$-SCLI framework (covering p=1p{=}1 one-step methods) with matching bounds; and Arjevani and Shamir [2] extended dimension-free iteration lower bounds to finite-sum settings, covering variance-reduced families. Complementary results for randomized first-order algorithms appear in Woodworth and Srebro [36]. More recently, Carmon, Duchi, Hinder, and Sidford [7] established sharp lower bounds for finding first-order stationary points in smooth convex and nonconvex optimization, clarifying the fundamental $1/T$ (convex) and $1/\sqrt{T}$ (nonconvex) barriers in gradient-norm complexity under standard first-order oracles. A comprehensive and recent treatment of finite-sum lower bounds is given by Han, Xie, and Zhang [21].
Multiobjective optimization algorithms
@@
Classical foundations and broad context are covered in the monograph by Miettinen [25] and the survey of Marler and Arora [24]. For descent-type methods, the steepest–descent family beginning with [15] and including cone-ordered variants and refined convergence analyses [11, 9] provides global convergence to Pareto-stationary points. Worst-case complexity guarantees of order $\mathcal{O}(\varepsilon^{-2})$ for trust-region and regularization frameworks in unconstrained MOO were established in [20]. Complexity aspects for gradient descent in smooth MOO are analyzed in [13]. Universal nonmonotone line-search frameworks for non-convex MOO with convex constraints have also been proposed [31]. When progress is tracked through standard merit functions or stationarity surrogates, the resulting upper bounds are typically sublinear (e.g., $\mathcal{O}(1/k)$) in smooth convex settings [35, 18]. Second-order frameworks — Newton-type and SQP — offer globalization with fast local behavior: globally convergent Newton methods for MOO with fast (often superlinear) local rates under standard regularity are developed in [19], while SQP-based schemes for constrained problems are given in [14]. Conditional-gradient (Frank–Wolfe) methods have been developed for vector/multiobjective optimization with global convergence to weakly efficient (Pareto-stationary) solutions under cone-convexity assumptions [8]. Away-step and related variants can accelerate convergence and, under additional geometric or curvature assumptions, deliver faster (sometimes linear) rates [17, 18], with adaptive generalized conditional-gradient methods extending this line [16]. Accelerated first-order schemes tailored to the multiobjective setting continue to mature: an accelerated proximal-gradient method with $\mathcal{O}(1/k^{2})$ decrease for an appropriate merit measure is established in [34], and recent work clarifies monotonicity properties of multiobjective APG [29]. Complementary practical gradient-based schemes and stationarity criteria, such as multiple-gradient descent, are surveyed in [12], and links to multi-task learning further motivate stationarity-based goals and algorithm design [33].
Our position
While scalarization-based upper bounds are classical, a systematic oracle-complexity theory for Pareto stationarity that (i) distinguishes oblivious one-step from both oblivious span methods and general adaptive first-order methods, (ii) provides robust non-degenerate hard MOO instances, and (iii) connects polynomial-degree lower bounds to geometric stationary-point lower bounds, has been missing. We supply such a framework and import sharp scalar stationary-point lower bounds into MOO via an explicit lifting reduction.
@@ 3.1 Notation
For $m\geq 2$ objectives $F(x)=(f_{1}(x),\dots,f_{m}(x))$, the unit simplex is
$\Delta^{m}:=\{\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\}.$
For $\lambda\in\Delta^{m}$, the scalarization is $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ and $x^{\star}_{\lambda}\in\arg\min_{x}f_{\lambda}(x)$ denotes an arbitrary minimizer.
In convex settings we use the scalarization value gap at initialization
$\Delta_{\lambda}(x^{(0)}):=f_{\lambda}(x^{(0)})-\inf_{x}f_{\lambda}(x).$
When the Pareto set $\mathcal{P}$ is well-defined (convex case; see Proposition 3.8), we write
$R:=\mathrm{dist}(x^{(0)},\mathcal{P}).$
In the nonconvex extension we will instead refer to the set of Pareto-stationary points $\mathcal{S}:=\{x:\mathcal{G}(x)=0\}$ when needed.
Iterates are $x^{(t)}$ for $t\in\mathbb{N}_{0}$. $T\in\mathbb{N}$ is the number of oracle calls/iterations.
The condition number is $\kappa:=L/\mu$ when strong convexity applies. We occasionally write $\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}$ in spectral formulas.
For quadratics $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, we use the spectral variable $\zeta$ and real polynomials $p_{t}$ (with $\deg p_{t}\leq t$, $p_{t}(0)=1$) to express errors via $x^{(t)}-x^{\star}=p_{t}(H)\bigl(x^{(0)}-x^{\star}\bigr)$. Chebyshev polynomials of the first kind are denoted $T_{t}(\cdot)$.
Standard Landau symbols $O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ are used; $\log$ denotes the natural logarithm.
For nonnegative quantities $A$ and $B$, we write $A\asymp B$ (equivalence up to universal constants) if there exist absolute constants $c_{1},c_{2}>0$ (independent of problem parameters such as $L,\Delta,R,T,m,d$) such that $c_{1}B\leq A\leq c_{2}B$. We use this notation only when the implicit constants are universal; otherwise we state explicitly which parameters they may depend on.
The vectors $(e_{i})_{i=1}^{m}$ are the canonical basis of $\mathbb{R}^{m}$.
@@ 3.2 Problem formulation
We consider unconstrained multiobjective problems
@@
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable. Convexity and strong convexity will be imposed in the relevant sections.
@@ 3.3 Pareto optimality and first-order conditions
Relation to Pareto merit functions
Besides stationarity gaps, the MOO literature also uses merit (value-gap) functions. Assume that each $f_{i}$ is convex and differentiable. For instance, consider the Pareto merit function
$u_{0}(x)\;:=\;\sup_{z\in\mathbb{R}^{d}}\min_{i=1,\dots,m}\bigl(f_{i}(x)-f_{i}(z)\bigr),$
which vanishes at weakly Pareto optima and is positive otherwise (see, e.g., [34]).
Let $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ for $\lambda\in\Delta^{m}$, and assume $\arg\min f_{\lambda}\neq\emptyset$ for all $\lambda\in\Delta^{m}$; pick any minimizer $x_{\lambda}^{\star}\in\arg\min f_{\lambda}$ and write $f_{\lambda}^{\star}:=f_{\lambda}(x_{\lambda}^{\star})$. We can upper bound $u_{0}(x)$ by scalarization gaps in three simple steps. Fix $x$ and $z$ and apply the identity $\min_{i}a_{i}=\min_{\lambda\in\Delta^{m}}\sum_{i}\lambda_{i}a_{i}$ with $a_{i}=f_{i}(x)-f_{i}(z)$, to obtain
$\min_{i=1,\dots,m}\bigl(f_{i}(x)-f_{i}(z)\bigr)=\min_{\lambda\in\Delta^{m}}\sum_{i=1}^{m}\lambda_{i}\bigl(f_{i}(x)-f_{i}(z)\bigr)=\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Substituting into the definition of $u_{0}$ yields
$u_{0}(x)=\sup_{z\in\mathbb{R}^{d}}\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Next, we use the elementary min–max inequality
$\sup_{z}\min_{\lambda}\Phi(z,\lambda)\leq\min_{\lambda}\sup_{z}\Phi(z,\lambda)$
with $\Phi(z,\lambda)=f_{\lambda}(x)-f_{\lambda}(z)$ to obtain
$u_{0}(x)\;\leq\;\min_{\lambda\in\Delta^{m}}\sup_{z\in\mathbb{R}^{d}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Finally, for each fixed $\lambda$, since $f_{\lambda}(x)$ does not depend on $z$,
$\sup_{z\in\mathbb{R}^{d}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr)=f_{\lambda}(x)-\inf_{z\in\mathbb{R}^{d}}f_{\lambda}(z)=f_{\lambda}(x)-f_{\lambda}^{\star},$
and therefore
$u_{0}(x)\;\leq\;\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}^{\star}\bigr).$
By convexity, for each $\lambda$,
$f_{\lambda}(x)-f_{\lambda}^{\star}\leq\langle\nabla f_{\lambda}(x),\,x-x_{\lambda}^{\star}\rangle\leq\|\nabla f_{\lambda}(x)\|\,\|x-x_{\lambda}^{\star}\|.$
Therefore,
$u_{0}(x)\;\leq\;\bar{R}(x)\,\mathcal{G}(x),\qquad\bar{R}(x):=\sup_{\lambda\in\Delta^{m}}\|x-x_{\lambda}^{\star}\|.$
Moreover, if each $f_{i}$ is $\mu$-strongly convex, then so is every $f_{\lambda}$, and the Polyak–Łojasiewicz inequality yields $f_{\lambda}(x)-f_{\lambda}^{\star}\leq\frac{1}{2\mu}\|\nabla f_{\lambda}(x)\|^{2}$; hence
$u_{0}(x)\;\leq\;\frac{1}{2\mu}\,\mathcal{G}(x)^{2}.$
In particular, under a boundedness condition ensuring $\bar{R}(x)<\infty$ on the region of interest (or under strong convexity), any convergence guarantee stated in terms of the stationarity gap $\mathcal{G}(x)$ immediately implies a corresponding bound for the merit function $u_{0}(x)$.
Equivalence of stationarity notions
@@
    (Convex-hull stationarity) There exists $\lambda\in\Delta^{m}$ such that
    $\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)=0$
    (equivalently, $\mathcal{G}(x)=0$).
@@
Necessary and sufficient conditions